In a series RL circuit, L = 4H, R = 100 ohms and E = 200V . Find the values of current as a function of time. Assume that the initial current is zero. Find the current when t = 2 secons.

1
Expert's answer
2022-03-08T15:12:01-0500

Solution;

Using the equation;

"L\frac{di}{dt}+Ri=E"

Rewrite;

"\frac{Di}{dt}+\frac{R}{L}I=\frac EL"

The integrating factor;

"I.F=e^{\int \frac RLdt}=e^{\frac{Rt}{L}}"

I put into the equation;

"e^{\frac{RT}{L}}[\frac{di}{dt}+\frac RLi]=e^{\frac{Rt}{L}}\frac EL"

Integrating respectively;

"I=\frac ER+ce^{\frac{-Rt}{L}}"

At t=0,i=0;

"c=-\frac ER"

"i=\frac ER-\frac ERe^{\frac{-Rt}{L}}"

Substitute the values;

"i=\frac{200}{100}-\frac{200}{100}e^{\frac{-100×2}{4}}"

"I=9.837A"